OpenStudy (anonymous):

What are the next three terms for the sequence 7, 10, 13, 16, 19, …?

OpenStudy (anonymous):

10 = 7+3 13 = 10 + 3 16 = 13 + 3 19 = 16 +3 Do you see it?